Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 377     Entry time: 30 Sep 2023, 23:08
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 6.3 keV, 3-4 mJ/pulse, 1-352 pulses at 1.13 MHz

Optical laser delivery

  • PP laser, 1030 nm, 352 pulses at 1.1 MHz, stable delivery
  • ns laser for rayleigh scattering, stabe delivery

Achievements/Observations

  • Several tens of datasets were collected using two different laser polarization (0 degree and 45 degree)
  • During the data collections PP laser phase shifter delay scans were used.
  • PP laser time aligned to zero and spatial overlapped was checked.

Issues

  • The cross talks between SASE-1 and SASE-2 was more than it expected.
  • Beam position, pulse energy as well as intra-train pulse energy were varrying. (BKR managed to solved it before the evening)
ELOG V3.1.4-7c3fd00